# Blinded by Generated Contexts: How Language Models Merge Generated and Retrieved Contexts for Open-Domain QA?

### v4

@@ A.2.1 Length Distribution across LLMs
Generate a background context from Wikipedia to answer the given question {#question} Keep the length of the document around $n$ words
@@
As a result, Figure 9 shows the length distribution of retrieved contexts and contexts generated by various LLMs. The length distribution of retrieved contexts is more concentrated as they consist of text limited to precisely 100 words, along with their titles (Karpukhin et al., 2020). The variation in the length of different retrieved contexts is solely due to the differences in title lengths.
A.2.2 Length Distribution with Different Control Methods
In Section 5.3, we employ three methods, “Nature”, “Trunc.” and “S-Trunc.”, to vary the completeness of generated contexts, while controlling the length at the same time. Figure 10 illustrates the length distribution for generated contexts corresponding to these methods. From the results, we can observe that the contexts generated by original GenRead (Yu et al., 2022) are significantly longer compared to the retrieved contexts.
Table 8 illustrates the average similarity and completeness of three types of generated contexts. “Nature”, “Trunc.” and “S-Trunc.” result in contexts with similar average length and similarity, with preliminary differences in completeness.
Context Completeness NQ-AIR TQA-AIR Sentence Semantic Length Jaccard Length Jaccard GPT-3.5 Nature 106.0 0.187 105.0 0.398 S-Trunc 104.9 0.184 106.3 0.397 Trunc 107.4 0.187 103.6 0.397 Llama2-13b Nature 109.7 0.202 105.5 0.313 S-Trunc 105.9 0.193 104.8 0.293 Trunc 107.4 0.196 106.3 0.295
Table 8: Average length and similarity of generated context with different control methods. Three types of methods create contexts with comparable average length and similarity. The core difference lies in the completeness.
@@ A.2.2 Length Distribution with Different Control Methods
Figure 10: Length distribution of generated and retrieved contexts on the NQ dataset with GPT-3.5 as the generator. "genread" represents the contexts generated by the original GenRead method (Yu et al., 2022). "trunc", "strunc", and "nature" are the generated contexts using three different methods to control the length.
A.3 Dataset Size
@@
A.4 Effect of Context Order
@@
Table 10: $\mathrm{DiffGR}$ with different context order on NQ-AIR and TQA-AIR datasets. GPT-3.5 serves as the generator and reader.
@@
A.5 More Results on AIG Datasets
A.6 More Details about Confirmation Bias
A.6.1 Experiment Setting
We construct counter-memory contexts for each instance on the original AIR context-conflicting subsets, as outlined below:
Counter-Memory Answers Preparation. For each question on the AIR subsets, we substitute the original memory answer (e.g., “Canada”) with a same-type yet distinct entity (e.g., “United States”), which serves as the counter-memory answer. Concretely, we employ ChatGPT to associate a different entity of the same type. It is noteworthy that the prompt does not incorporate the question, ensuring that the LLM’s counter-memory answer is merely a categorical equivalent rather than a memory related to the question.
Give you a reference word, transform it into an analogous word.
Reference: Missouri River
Analogous Word: Mississippi River
Reference:
{memory answer}
Analogous Word:
Furthermore, to facilitate the calculation of $\mathrm{DiffGR}$, we only retain those samples where the counter-memory answer also diverges from $a_{\phi}^{\gamma}$, the answer provided by the retrieved context. Formally, the counter-memory answer satisfies $a_{\phi}^{\varrho^{\prime}}\neq a_{\phi}^{\varrho}$ and $a_{\phi}^{\varrho^{\prime}}\neq a_{\phi}^{\gamma}$.
Counter-Memory Contexts Generation. We enforce LLMs to generate a counter-memory context that supports the counter-memory answer. This approach is inspired by the success of using LLMs to generate misinformation (Pan et al., 2023).
Generate a background document in support of the given opinion to the question. Keep the length of the document around $n$ words. Question: {question} Opinion: {counter memory answer} Document:
Following Section 3, the prompt also incorporates a length constraint to mitigate the influence of length, as evidenced by the length statistics presented in Table 11.
Answer Consistency Checking. To verify that the counter-memory context actually supports the counter-memory answer, we retain only those instances where the predicted answer, derived exclusively from the counter-memory context, exactly matches the counter-memory answer:
$a_{\phi}^{\varrho^{\prime}}=\phi(q,d_{1}^{\varrho^{\prime}})$
Context NQ-AIR TQA-AIR ConflictQA(PopQA) Length Jaccard Length Jaccard Length Jaccard Retrieved 107.5 0.111 106.3 0.184 - - GPT-3.5 Gen 105.9 0.194 104.9 0.385 62.14 0.169 Ctr 105.7 0.251 103.8 0.499 100.4 0.236 Llama2-13b Gen 107.4 0.197 104.3 0.319 - - Ctr 106.7 0.165 103.5 0.310 - -
Table 11: Average length and Jaccard similarity of different contexts. Gen, Ret, and Ctr respectively represent generated contexts, retrieved contexts, and counter-memory contexts. “-” means the context is not included in this dataset. Detailed discussion about Jaccard similarity is shown in Section 5.2.
Reader & Generator NQ-AIR TQA-AIR
GPT-3.5 504 596
Llama2-13b 883 1376
Table 12: The size of context-conflicting subsets with counter-memory contexts.
Following the above procedures, we obtain subsets from original AIR datasets, with each instance encompassing a question $q$, a generated context $d_{1}^{\varrho}$, a retrieved context $d_{1}^{\gamma}$, a counter-memory context $d_{1}^{\varrho^{\prime}}$ and the associated answers for these contexts. Table 12 shows the size of developed context-conflicting subsets with counter-memory contexts.
A.6.2 Discussion
In Table 4, we observe a phenomenon: GPT-3.5 exhibits a stronger bias towards counter-memory contexts compared to the original generated contexts, which does not exist in Llama2-13b. This phenomenon is also observed in Xie et al. (2023) (Table 5 in their work). Upon further investigation, we discover that this may be due to the relatively higher similarity between the counter-memory context generated by GPT-3.5 and the question, as shown in Table 11. In contrast, the generated context and counter-memory context for Llama2-13b exhibit approximately equal levels of similarity. Furthermore, we analyze the original ConflictQA (PopQA) datasets introduced by Xie et al. (2023) and identify a similar disparity in context similarities, as shown in Table 11. This means that previous works may ignore some significant factors to LLMs’ preference, such as similarity (investigated in Section 5.2).
A.7 More Details and Results about Similarity
A.7.1 Similarity Metric
@@
We employ Jaccard similarity to assess the term-based overlap, and BERTScore (Zhang et al., 2020) for evaluating the semantic similarity between contexts and questions. To mitigate the effect of length discrepancies between contexts and questions, we calculate the similarity at the sentence level and then aggregate them to derive the overall context-question similarity. In this work, we adopt a maximum aggregation strategy due to the single-hop nature of the NQ and TQA datasets, where the majority of questions can be answered using a small subset of sentences. We also try the average aggregation strategy and observe similar results.
@@
A.7.2 Similarity Distribution
Figure 13 and 14 show the similarity distribution of retrieved and generated contexts across various generators. All LLM-generated contexts exhibit a higher similarity over retrieved contexts.
A.7.3 Effect of Similarity.
Figure 15 demonstrates a general trend that on slices with a smaller average similarity gap, LLMs exhibit a reduced preference for generated context.
@@ A.7.2 Similarity Distribution
Figure 15: The generation ratio in slices with different average $\Delta\operatorname{sim}$. $\Delta\operatorname{sim}$ is the difference in similarity between the generated context and the retrieved context.
A.7.4 Cases about Similarity
Table 13 shows examples that contain contexts with different similarities to the question. The contexts with high similarity typically directly support answering by repeating the phrasing in the question. Conversely, the contexts with low similarity introduce more challenges, often necessitating an understanding of synonyms and even some inferences. These observations indicate that text similarity can partly reflect the relevance between a question and a context, as well as the difficulty the LLM encounters in identifying potential answers.
@@ A.7.4 Cases about Similarity
A.8 Cases about completeness

### v6

@@ A.1 Length Control for Generated Contexts
Generate a background context from Wikipedia to answer the given question {#question}. Keep the length of the document around $n$ words.
@@
As a result, Table 8 illustrates the average length of different contexts. Figure 9 shows the length distribution of retrieved contexts and contexts generated by various LLMs. The length distribution of retrieved contexts is more concentrated as they consist of text limited to precisely 100 words, along with their titles (Karpukhin et al., 2020). The variation in the length of different retrieved contexts is solely due to the differences in title lengths.
@@
A.2 Dataset Size
@@
Appendix B Additional Results
B.1 More Results on AIG Datasets
B.2 More Results about Effect of Parametric Knowledge
Table 10 illustrates the size of the filtered AIR datasets with the constraint “$a_{\phi}^{\text{llm}}\neq a_{\phi}^{\gamma}\neq a_{\phi}^{\varrho}$”. GPT-4 corresponds to fewer samples because it has better parametric knowledge, i.e., $a_{\phi}^{\text{llm}}$ is more likely to be correct. Table 11 illustrates the proportion of LLMs choosing the answer provided by generated contexts, retrieved contexts, or parametric knowledge.
Reader Generator NQ (12367) TQA (20150) NQ-AIG NQ-AIR TQA-AIG TQA-AIR GPT 4 GPT 4 254 342 259 180 GPT 4 GPT 3.5 220 381 - - GPT 4 Llama2-13b 160 527 - - GPT 4 Llama2-7b 124 545 - - GPT 3.5 GPT 4 653 423 796 237 GPT 3.5 GPT 3.5 500 457 524 322 GPT 3.5 Llama2-13b 271 665 359 574 GPT 3.5 Llama2-7b 211 708 279 696 Llama2-13b GPT 4 1563 478 2321 309 Llama2-13b GPT 3.5 1318 553 1889 467 Llama2-13b Llama2-13b 633 841 928 1020 Llama2-13b Llama2-7b 515 1018 750 1358 Llama2-7b GPT 4 1896 479 3163 418 Llama2-7b GPT 3.5 1682 557 2746 591 Llama2-7b Llama2-13b 908 884 1681 1238 Llama2-7b Llama2-7b 753 1002 1272 1538
Table 10: The number of data in selected subsets where $a_{\phi}^{\text{llm}}\neq a_{\phi}^{\gamma}\neq a_{\phi}^{\varrho}$.
Reader Generator NQ-AIR TQA-AIR $\rho_{\text{gen}}$ $\rho_{\text{ret}}$ $\rho_{\text{llm}}$ $\mathrm{DiffGR}$ $\rho_{\text{gen}}$ $\rho_{\text{ret}}$ $\rho_{\text{llm}}$ $\mathrm{DiffGR}$ GPT 4 GPT 4 66.08 18.71 2.05 0.5586 76.67 13.89 2.78 0.6933 GPT 4 GPT 3.5 66.40 23.62 1.84 0.4752 - - - - GPT 4 Llama2-13b 58.63 28.84 2.85 0.3406 - - - - GPT 4 Llama2-7b 62.20 26.42 2.39 0.4037 - - - - GPT 3.5 GPT 4 65.25 16.08 0.95 0.6047 68.78 17.72 3.38 0.5902 GPT 3.5 GPT 3.5 67.83 12.91 0.88 0.6802 72.05 16.15 1.55 0.6338 GPT 3.5 Llama2-13b 67.22 15.34 1.20 0.6284 70.21 16.38 2.61 0.6217 GPT 3.5 Llama2-7b 64.55 16.24 1.13 0.5979 74.14 13.65 1.87 0.6890 Llama2-13b GPT 4 61.72 25.73 2.51 0.4115 57.28 33.01 1.29 0.2688 Llama2-13b GPT 3.5 62.39 26.76 1.08 0.3996 61.88 26.34 1.71 0.4029 Llama2-13b Llama2-13b 69.92 18.67 1.43 0.5785 72.55 17.75 1.96 0.6069 Llama2-13b Llama2-7b 69.16 18.96 1.87 0.5697 75.04 15.54 1.62 0.6569 Llama2-7b GPT 4 54.28 24.43 1.46 0.3793 47.13 33.97 3.59 0.1622 Llama2-7b GPT 3.5 52.24 27.29 1.80 0.3138 54.31 28.76 2.20 0.3075 Llama2-7b Llama2-13b 61.65 19.46 3.17 0.5202 64.46 20.76 1.86 0.5128 Llama2-7b Llama2-7b 60.38 21.16 2.20 0.4810 66.25 18.40 3.12 0.5653
Table 11: The columns labeled $\rho_{\text{gen}}$, $\rho_{\text{ret}}$, and $\rho_{\text{llm}}$ respectively represent the proportion (%) of responses that match the answer provided by generated contexts, retrieved contexts, and internal parametric knowledge of LLMs.
B.3 Effect of Context Order
@@
Table 12: $\mathrm{DiffGR}$ with different context order on NQ-AIR and TQA-AIR datasets. GPT-3.5 serves as both the generator and reader.
@@
NQ-AIR TQA-AIR Llama2-13b 632 702 GPT-3.5 323 181
Table 13: The number of samples filtered with a BERTScore difference of less than 0.05.
B.4 More Results about Effect of Context Completeness
(a) Length Distribution. In §5.3, we employ three methods, “Nature”, “Trunc.” and “S-Trunc.”, to vary the completeness of generated contexts, while controlling the length at the same time. Figure 11 illustrates the length distribution for generated contexts corresponding to these methods. From the results, we can observe that the contexts generated by original GenRead (Yu et al., 2022) are significantly longer compared to the retrieved contexts.
(b) Similarity and Length Control. Table 13 presents the data quantity after filtering out samples where three types of generated contexts exhibit significant differences in similarity. Table 14 illustrates the average similarity and completeness of three types of generated contexts. “Nature”, “Trunc.” and “S-Trunc.” result in contexts with Comparable average length and similarity, with preliminary differences in completeness.
Context Completeness NQ-AIR TQA-AIR Sentence Semantic Length Jaccard BERTScore Length Jaccard BERTScore Llama2-13b Retrieved 107.8 0.1157 0.8534 106.6 0.1832 0.8630 Nature 109.4 0.1859 0.8755 105.0 0.3178 0.8909 S-Trunc 106.1 0.1734 0.8719 105.0 0.2868 0.8833 Trunc 107.8 0.1765 0.8726 106.6 0.2890 0.8837 GPT-3.5 Retrieved 107.1 0.1162 0.8552 106.3 0.1836 0.8622 Nature 105.7 0.1748 0.8730 105.0 0.3993 0.9044 S-Trunc 105.8 0.1733 0.8730 104.8 0.4024 0.9044 Trunc 107.1 0.1769 0.8736 106.3 0.4027 0.9044
Table 14: Average length and similarity of contexts for GPT-3.5 and Llama2-13b models. The three types of generated contexts exhibit similar average lengths and similarity, with the primary distinction being their completeness.
Figure 11: Length distribution of generated and retrieved contexts on the NQ dataset with GPT-3.5 as the generator. “genread” represents the contexts generated by the original GenRead method (Yu et al., 2022). “trunc”, “strunc”, and “nature” are the generated contexts using three different methods to control the length.
B.4.1 Cases about Completeness
B.5 More Results about Effect of Text Similarity
@@
(a) Similarity Metric. We employ Jaccard similarity to assess the term-based overlap, and BERTScore (Zhang et al., 2020) for evaluating the semantic similarity between contexts and questions. To mitigate the effect of length discrepancies between contexts and questions, we calculate the similarity at the sentence level and then aggregate them to derive the overall context-question similarity. In this work, we adopt a maximum aggregation strategy due to the single-hop nature of the NQ and TQA datasets, where the majority of questions can be answered using a small subset of sentences. We also try the average aggregation strategy and observe similar results.
@@
(b) Similarity Distribution. Figure 13 and 14 show the similarity distribution of retrieved and generated contexts across various generators. All LLM-generated contexts exhibit a higher similarity over retrieved contexts.
(c) Effect of Similarity. Figure 15 demonstrates a general trend that on slices with a smaller average similarity gap, LLMs exhibit a reduced preference for generated context.
@@
Figure 15: The generation ratio in slices with different average $\Delta\operatorname{sim}$. $\Delta\operatorname{sim}$ is the difference in similarity between the generated context and the retrieved context. The LLM marked in the figure serves both as a generator and a reader.
(d) Cases about Similarity. Table 16 shows examples that contain contexts with different similarities to the question. The contexts with high similarity typically directly support answering by repeating the phrasing in the question. Conversely, the contexts with low similarity introduce more challenges, often necessitating an understanding of synonyms and even some inferences. These observations indicate that text similarity can partly reflect the relevance between a question and a context, as well as the difficulty the LLM encounters in identifying potential answers.
@@
B.6 The Effectiveness of Exactly Matching
Table 17 demonstrates that the proportion of “Others” is significantly lower relative to the disparities between “Gen” and “Ret”, its impact on the conclusions of this paper is negligible.